OpenStudy (misssmartiez):

SIMPLIFY: (2^5 / 3^2)^4 I think its A. (I will draw it out.) A) 2^20/3^8 B) -2^9/-3^8 C) 8^5/12^2 D) 2/3^2

TheSmartOne (thesmartone):

Correct (:

OpenStudy (misssmartiez):

Thanks!

OpenStudy (misssmartiez):

So I simply multiply the exponents?

TheSmartOne (thesmartone):

Anytime! This question applies two rules: \(\Large (\frac{a}{b})^m = \frac{a^m}{b^m}\) \(\Large (a^b)^c = a^{bc}\)

OpenStudy (misssmartiez):

Ok, thanks!

TheSmartOne (thesmartone):

Anytime! :)